A current of 2 A flowing through a coil of 100 turns gives rise to a magnetic flux of 5 × 10−5 Wb per turn. Calculate the magnetic energy associated with the coil


1
Expert's answer
2022-01-14T09:49:14-0500

"W=\frac{n\Phi I}2=\frac{100\cdot 5\cdot 10^{-5}\cdot 2}2=5~mJ."